13. A die is thrown once. Find the probability of getting

(i) a prime number;

(ii) a number lying between 2 and 6;

(iii) an odd number.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Total possible events when a dice is thrown = 6 (1, 2, 3, 4, 5, and 6) P(E) = (Number of favourable outcomes/ Total number of outcomes) (i) Total number of prime numbers = 3 (2, 3 and 5) P (getting a prime number) = 3/6 = ½ = 0.5 (ii) Total numbers lying between 2 and 6 = 3 (3, 4 and 5) P (getting a number between 2 and 6) = 3/6 = ½ = 0.5 (iii) Total number of odd numbers = 3 (1, 3 and 5) P (getting an odd number) = 3/6 = ½ = 0.5